OpenStudy (abbycross167):

Can someone please help me fix this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression -

OpenStudy (abbycross167):

(4x^2-6x-18) (x^2+1) ------------ * ---------- (x+1) (x^2-6x+9)

OpenStudy (abbycross167):

these are the Steps I did- (4x^3-6x-18)(x^2+1) ---------------------- (x+1)(x^2-6x+9) (4x^2-6x-18) = 2(2x^2-3x-9)= 2(2x^2-6x+3x-9)= 2(2x^2-6x)+(3x-9)= 2(2x^2(x-3)+3(x-3) 2(x-3)(2x+3)= 2x^3 ------------- (x+1)(x^2-6x+9)

OpenStudy (abbycross167):

This is what my teacher said was wrong - Make your steps look like this example - (8y) (7x) ------ * ------ (14x^2) (8z^3) Steps-- 1 1 1 1 1 2*2*2*y * 7 * z' ---------------------- 2*7*x*x 2*2*2*z*z*z 1 1 1 1 1 1*1*1*y*1*1 ------------------- 2*1*1*x*1*1*1*z*z*z y 2xz3

OpenStudy (phi):

the first problem becomes \[ \frac{2(x-3)(2x+3)(x^2+1) }{(x+1) (x-3)^2}\] we can cancel (x-3) from the top and bottom. nothing else simplifies \[ \frac{2(2x+3)(x^2+1) }{(x+1) (x-3)}\]

OpenStudy (phi):

what is the question for the 2nd problem ?

OpenStudy (abbycross167):

You mean the example?

OpenStudy (phi):

oh, that was an example? never mind. are you ok with the first problem ?

OpenStudy (phi):

I see you factored (4x^2-6x-18) into 2(x-3)(2x+3) but I can't follow what you did after that ?

OpenStudy (phi):

the other step would be to factor (x^2-6x+9) (which is in the "bottom")

OpenStudy (abbycross167):

I think my teacher just wants me to rewrite the steps so they will make more sense, but I'm not sure how to rewrite them

OpenStudy (phi):

*** these are the Steps I did- (4x^3-6x-18)(x^2+1) ---------------------- (x+1)(x^2-6x+9) **** I assume you mean 4x^2 not 4x^3 , right ?

OpenStudy (phi):

can you factor (x^2-6x+9)

OpenStudy (abbycross167):

It's 4x^2 sorry I messed up typing it

OpenStudy (phi):

ok. meanwhile, can you factor x^2-6x+9 ?